OpenStudy (anonymous):

what is 9 divided by 4.50

OpenStudy (anonymous):

450 divided by 9 = 50

OpenStudy (anonymous):

\[9/4.50=2\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[9/4.50=2. \] \[Because 4.50*2=9 \]